OpenStudy (anonymous):

Marcus spent 10 hours doing his homework last week. This week he spent 11 hours doing homework. He says that he spent 110% more time doing homework this week. Is he correct? Show your work to justify your decision. Giving medal to anyone who answers this. I know this is history but please help me with this.

OpenStudy (e.mccormick):

OK, so if you have something that is a percentage of something else, do you know how to set that up as a formula?

OpenStudy (anonymous):

This isn't correct. Technically speaking, you cannot go over 100% because it is the maximum of what you can go to.

OpenStudy (e.mccormick):

@FlameGod You do not understand percentages as completely as you think. You can have more than 100%. For example, if I sell something for 3 times as much as it costs, have 200% profit.

OpenStudy (e.mccormick):

@ashley_13 Let me show an few examples. 5 is, 50% of 10. So, what formulas can I make with that: \(10 \times 50\% = 5\) Now, 50% means the same as \(\dfrac{50}{100}\) which becomes .5 as a decimal, so I could have written that as: \(10 \times \dfrac{50}{100} = 5\) or \(10 \times .5 = 5\) OK, but what if I am missing one of them? \(10 \times 50\% = ?\) Simplest case, I just multiply. \(10 \times ?\% = 5\) \(? \times 50\% = 5\) Those would need some more work. A little algebra to move things where they need to be to find the unknown.

OpenStudy (anonymous):

@e.mccormick i dont fully understand?

OpenStudy (e.mccormick):

Well, they are asking you to confirm if this statment is true or not. So, you need to test it mathematically.

OpenStudy (anonymous):

all i got is 10x110%=11 so do i put like 110 over 100 or something? @e.mccormick

OpenStudy (e.mccormick):

Whichever way you want. There are three things. Remove one then solve the other two to see if you get the same thing. Also, I just noticed you posted this in History... Should have aske it in the math area. =)

OpenStudy (anonymous):

i realize tht but i culdnt figure out how to get to the math

OpenStudy (e.mccormick):

If you solve any of these: \(10\times 110\%=?\) \(10\times ?\%=11\) \(?\times 110\%=11\) And you get the original, it is true. Up at the top, Find More Subjects.

OpenStudy (anonymous):

true i tink

OpenStudy (anonymous):

@e.mccormick i multiplied 10 by 110% and got ll so it it true?

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

ok so the answer is yes he is correct because i multiplied 10 by 110% and got 11 ? @e.mccormick

OpenStudy (e.mccormick):

Yes, and the math you did is the justification.

OpenStudy (anonymous):

ok thank you so much you helped me a whole lot! :) @e.mccormick

OpenStudy (e.mccormick):

np. It is better to know how it works from an eample, then you can solve it... which you did.